OpenStudy (anonymous):

how do i find the y intercept of a parabola such as y=x^2+2x

OpenStudy (anonymous):

make x=0

OpenStudy (anonymous):

c =0

OpenStudy (anonymous):

? you were no help

OpenStudy (anonymous):

any equasion?

OpenStudy (anonymous):

x=0 to find y-intercept y=0^2+2(0) y-intercept at (0,0)

OpenStudy (anonymous):

is that always the case where u plug in 0 for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so there is a point on the parabola at 0,0

OpenStudy (anonymous):

sure is

OpenStudy (anonymous):

that's the vertex luis, quite different from the y-intercept

OpenStudy (anonymous):

what would be the axis of symetry for this problem?

OpenStudy (anonymous):

y=x^2+2x h = -2/2 = -1 ->k = (-1)^2 - 2 = -1

OpenStudy (anonymous):

what? any formula?

OpenStudy (anonymous):

-b/(2a) = -2/2=-1, x=-1, or you can complete the square as luis did. the axis of symmetry is the vertical line that goes through the x-coordinate of the vertex

OpenStudy (anonymous):

haha thanks though im a little confused about the vertex

OpenStudy (anonymous):

or highest if its maximum right?

OpenStudy (anonymous):

it will be a maximum if a<0 in ax^2+bx+c

OpenStudy (anonymous):